OpenStudy (anonymous):

4x – 3y < –12 x – 6y >_ 6 can you please show me what the graph would look like

OpenStudy (anonymous):

Can you show me how to graph x-6y greater than or equal to -18

OpenStudy (anonymous):

FIRST, change it into Slope-Intercept form ~ \[x - 6y \ge -18 \]Subtract x from both sides. \[-6y \ge -x - 18\]Divide both sides by -6. Flip the sign to <=. \[y \le \frac{ 1 }{ 6 } x + 3\]------------------------ Remember: y = mx + b m is slope b is y-intercept \[y \le \frac{ 1 }{ 6 } x + 3\]Slope is 1/6 y-intercept is 3 Graph the equation. Step 1) Plot the y-intercept (3) at point (0, 3). Step 2) Remember that:\[slope = \frac{ rise }{ run } = \frac{ 1 }{ 6 }\]So from point (0, 3) RISE 1 space up, and RUN 6 spaces to the right. Plot this point.

OpenStudy (anonymous):

And draw a line between the two points.

OpenStudy (anonymous):



OpenStudy (anonymous):